OpenStudy (anonymous):

find the second derivative of f(x)= xe^(x/4) find the second derivative of f(x)= xe^(x/4) @Mathematics

OpenStudy (anonymous):

\[f(x)=xe ^{x/4}\]\[f \prime(x)=1/4*e ^{x}(x+1)\]\[f \prime \prime (x)=1/4*e ^{x}(x+2)\]

OpenStudy (anonymous):

Err, for the second derivative I meant:\[f \prime \prime(x)=1/16e ^{x/4}*(x+8)\]